PHOTOCHEMISTRY OF HALOGENATED TRANSITION METAL DIANIONS

\underline{ALEXANDER N. TARNOVSKY}, IGOR L. ZHELDAKOV, EVGENIIA V. BUTAEVA, and AN-DREY S. MERESHCHENKO, Department of Chemistry, Bowling Green State University, Bowling Green, OH, 43402.

Separation of two negative charges in aqueous PtBr_{6}^{2−} is investigated by means of femtosecond broadband pumpprobe spectroscopy using several excitation wavelengths, in concert with time-resolved x-ray absorption spectroscopy and DFT/TDDFT calculations.